If y varies directly with x, and y = 54 when x = 9, find the constant of variation k.

Answer by ankor@dixie-net.com(22740)

If y varies directly with x, and y = 54 when x = 9, find the constant of variation k.
:
y = k*x
:
54 = 9k
:
54/9 = k
:
k = 6
:
Check:
y = 6*9
y = 54